# Beyond Position: the emergence of wavelet-like properties in Transformers

### v3

@@ 5 Experiments and Analysis
LLaMA 3.2 (1B) 32 1.333 9.980 0.983 0.850 0.568 0.019
Gemma 2 (2B) 8 1.809 8.103 0.986 0.866 0.225 0.028
Pythia (2.8B) 32 1.689 8.298 0.981 0.853 0.591 0.019
Qwen 2.5 (5B) 14 1.527 8.835 0.983 0.862 0.304 0.031
Mistral (7B) 32 2.217 6.729 0.983 0.850 0.657 0.014
LLaMA 3.1 (8B) 32 1.529 9.141 0.984 0.850 0.597 0.014
Pythia (12B) 40 2.006 6.462 0.984 0.850 0.597 0.014
Table 1: Comparative Analysis of Language Model Metrics
Model 16 tok. 32 tok. 64 tok.
LLaMA 3.2 (1B) 0.937 0.931 0.931
Gemma 2 (2B) 1.073 1.056 1.055
Pythia (2.8B) 0.942 0.940 0.940
Qwen 2.5 (5B) 1.106 1.103 1.103
Mistral (7B) 0.889 0.877 0.877
LLaMA 3.1 (8B) 0.878 0.876 0.877
Pythia (12B) 0.878 0.877 0.877
Table 2: Multi-Resolution Window Entropy Analysis
As we can see from Table 1 and Table 2, the remarkably consistent pattern across all models where correlation remains near-perfect (0.98) at $0.5x$ scale but degrades to 0.850.85~{}0.850.85 at $0.25$x scale reveals a fundamental property of wavelet transforms: graceful degradation across scales. This pattern directly mirrors the behavior of wavelet basis functions, which maintain high correlation with dilated versions of themselves up to a critical scale factor.
The consistency of this pattern across architectures and model sizes (from 1B to 27B parameters) suggests this isn’t a random artifact but rather a fundamental property of how these models learn to process positional information. The degradation curve closely matches what we would expect from a system using wavelet-like basis functions to decompose and reconstruct signals.
Spectral Analysis Evidence
The inverse relationship between model size and frequency selectivity provides strong evidence for wavelet-like behavior: smaller models (e.g., LLaMA 1B) show high frequency selectivity (9.980) and low spectral entropy (1.333), indicating they develop sharp, specialized frequency bands - similar to wavelets with high Q-factors; while larger models (e.g., Pythia 12B) show lower selectivity (6.462) and higher spectral entropy (2.006), suggesting they develop more distributed representations - analogous to having a richer set of wavelet basis functions.
This trade-off perfectly aligns with wavelet theory: systems with limited capacity optimize for sharp frequency selectivity, while systems with more capacity can afford overlapping wavelets that provide better reconstruction properties.
Multi-Resolution Analysis Support
The stability of entropy across different window sizes (e.g., Mistral 7B: [0.889, 0.877, 0.877]) provides crucial evidence for wavelet-like behavior. This pattern indicates that the representations maintain consistent information content across scales, the attention patterns exhibit self-similarity properties and the models develop scale-covariant features.
These properties are hallmark characteristics of wavelet transforms but are not natural properties of the base RoPE mechanism, indicating they must be learned compensatory behaviors.
Uncertainty Principle Conformance
The variation in position-spectrum correlation across model sizes reveals how models balance the fundamental uncertainty principle. In fact, smaller models (Gemma 2B: 0.224) show low correlation, indicating they maintain separate positional and frequency channels, while larger models (Mistral 7B: 0.657) show higher correlation, suggesting more integrated representations
This progression exactly matches what we would expect from a system evolving increasingly sophisticated wavelet-like properties: smaller models use simpler, more separated representations, while larger models develop more nuanced, integrated representations that better balance the position-frequency trade-off.
Frame Completeness Evidence
The systematic improvement in reconstruction error with model size (from 0.031 for Qwen 2.5 5B to 0.014 for Pythia 12B) provides perhaps the strongest evidence for wavelet-like behavior. This pattern shows that larger models develop more complete wavelet frames, that the representations become more orthogonal and efficient and the system learns to better approximate the completeness relation of wavelet frames.
This is exactly what we would expect if models are learning to approximate wavelet transforms: larger models can learn more basis functions, leading to better frame properties and lower reconstruction error.
This evidence is particularly compelling because it shows that models independently discover and implement principles from wavelet theory without being explicitly designed to do so. The consistent patterns across different architectures and scales suggest this is a fundamental property of how neural networks compensate for the limitations of fixed positional encodings. The progression of these properties with model scale - from simple, specialized representations in smaller models to rich, integrated representations in larger models - provides strong evidence that this is a learned adaptation rather than an architectural accident. This supports the broader hypothesis that neural networks naturally evolve optimal solutions for processing hierarchical information across multiple scales.
@@ 7 Implications
The practical implications of our findings are particularly compelling, understanding that attention heads naturally organize into frequency bands suggests new approaches to model initialization and architecture design. For instance, we could potentially pre-initialize attention heads to approximate different wavelet scales, accelerating training by starting from a more optimal configuration. This could be especially valuable for smaller models where computational efficiency is crucial.
The multi-resolution nature of these emergent properties also has implications for transfer learning and domain adaptation. Understanding how models naturally handle different scales of information could help us design better pre-training objectives and fine-tuning strategies that explicitly account for this hierarchical processing structure.
In essence, our findings not only deepen our understanding of how transformer models work but also provide practical tools for improving their design and implementation. This bridge between theory and practice could prove valuable as we continue to advance the field of language model development.
@@ 8 Conclusion
Our research into the relationship between rotary positional embeddings and attention patterns reveals a fascinating aspect of how large language models adapt to theoretical limitations. We have shown that transformer models naturally evolve wavelet-like properties, and that servers as a compensation mechanism for the inherent constraints of RoPE, with this behavior becoming more sophisticated as model scale increases.
The consistent pattern of frequency band distribution across different model scales, the systematic improvement in frame completeness with model size, and the remarkable stability of multi-resolution entropy all point to a learned adaptation that closely mirrors wavelet transform properties. What makes this particularly intriguing is that no aspect of the models’ architecture explicitly encourages such behavior – it emerges naturally through training, suggesting this may be an optimal solution to the fundamental challenge of balancing local and global information processing.
The progression of these properties with model scale is especially revealing. Smaller models develop simpler but more specialized frequency responses, while larger models evolve more nuanced, integrated representations. This scalability suggests that the wavelet-like behavior is not merely a coincidental feature but a fundamental property of how these models learn to process hierarchical information across multiple scales.
These findings have significant implications for future model development. Understanding that attention heads naturally evolve to approximate wavelet transforms could inform more efficient architectural designs, potentially leading to models that explicitly leverage this property rather than requiring it to be learned. This could be particularly valuable for smaller models, where computational efficiency is crucial.
Looking forward, these findings contribute to our understanding of how neural networks implement sophisticated mathematical principles, even when not explicitly designed to do so.
@@ 9 Limitations
Our study shown theoretically and with experiments that Transformer-based LLMs learn to process hierarchical information across multiple scales in a way that resemble a wavelet. However, limitations and questions remain to be addressed. First, our understanding of how these properties emerge during training is still limited: we shown that this behavior helps the model overcoming the theoretical limitations of RoPE, but we didn’t study whether is RoPE that induces the behavior itself. Second, our analysis studied this behavior at inference time, in fact we think that a possible future work would be studying the emergence of the wavelet-like patterns at training time.
An intriguing limitation we encountered involves the interaction between these wavelet-like properties and the model’s handling of ambiguous or context-dependent information. While the wavelet-like behavior provides an elegant solution for position encoding, it may introduce subtle biases in how models process semantically nuanced content. Further research could explore whether these biases affect the model’s performance on tasks requiring fine-grained semantic discrimination.
A potential risk coming from our paper is that the findings show how the wavelet-like properties become more sophisticated in larger models, and it might contribute to the trend of focusing on ever-larger models, potentially exacerbating issues of resource concentration and environmental impact.

### v4

@@ 5.3 Ablation Study
Llama-3.2-3B RoPE 24 2.425 $\pm$ 0.630 0.728 $\pm$ 0.349 0.038 $\pm$ 0.036 0.090 $\pm$ 0.040 -0.502 0.00012 $\pm$ 0.00008
flan-t5-base T5 12 2.696 $\pm$ 0.712 0.704 $\pm$ 0.480 0.627 $\pm$ 0.239 0.689 $\pm$ 0.231 -0.790 0.00004 $\pm$ 0.00003
bert-base BERT 12 2.449 $\pm$ 0.819 0.743 $\pm$ 0.446 0.507 $\pm$ 0.342 0.548 $\pm$ 0.350 -0.606 0.00007 $\pm$ 0.00006
gpt2 No PE 12 2.868 $\pm$ 0.767 0.514 $\pm$ 0.369 0.141 $\pm$ 0.164 0.152 $\pm$ 0.154 -0.672 0.00005 $\pm$ 0.00005
Table 4: Ablation study comparing different position encoding (PE) methods. We report mean and standard deviation across attention heads for key metrics.
To validate that the observed compensatory mechanisms are uniquely characteristic of RoPE’s implicit relative positioning, we conducted an ablation study comparing it against three alternative positional encoding schemes: T5’s relative position biases, BERT’s absolute positional embeddings, and a GPT-2 model with no explicit positional encoding (No PE). The results, detailed in Table 4, show that RoPE provides a uniquely effective solution for maintaining positional awareness across different scales.
The analysis highlights a clear hierarchy in scale invariance. RoPE stands out with an exceptionally low scale sensitivity ($S_{0.5}=0.038$), demonstrating a robust ability to preserve positional representations when sequence lengths change. At the other extreme, models with explicit, non-rotary encodings (T5 and BERT) proved to be comparatively rigid, with high scale sensitivity values of 0.627 and 0.507, respectively. T5’s rigidity is further underscored by its high spectral entropy (2.696) and the most strongly negative position-spectrum correlation ($\rho=-0.790$), indicating a significant conflict between positional focus and spectral organization. The GPT-2 model (No PE) presents a fascinating intermediate case. While it achieves moderate scale invariance ($S_{0.5}=0.141$), far superior to T5 and BERT, it is significantly less robust than RoPE. Interestingly, it accomplishes this not through sharp, focused attention, but through the opposite strategy. It registered the lowest frequency selectivity (0.514) and the highest spectral entropy (2.868), suggesting its attention patterns are spectrally diffuse. This implies that in the absence of positional guidance, the model adopts a high-entropy, less-structured approach that offers some resilience but lacks the precision of RoPE.
Ultimately, these findings underscore that RoPE’s architecture fosters wavelet-like properties that other positional encoding implementations don’t. It achieves scale-invariance that surpasses not only rigid, explicit encodings but also the more diffuse, emergent strategies of models with no positional guidance, marking it as a uniquely effective and well-balanced method for encoding position.
5.4 Evolution of wavelet-like features during training
To understand how this wavalet-like compensatory strategies develop, we analyzed Pythia 6.9b at different stages of its training, from initialization (step 0) to 143,000 steps. The results, summarized in Table 5, reveal that the model does not learn its representations linearly but instead undergoes distinct developmental phases.
Initially, at step 0 and step 128, the untrained model exhibits a simple default state: high frequency selectivity ($\sim 0.76$) and low spectral entropy ($\sim 2.29$). This indicates that the attention heads are undifferentiated, focusing on very basic, low-complexity patterns before significant learning has occurred.
A dramatic "exploratory phase" begins around step 512 and peaks precisely at step 1000. In this critical period, frequency selectivity plummets to its absolute minimum of 0.230, while spectral entropy surges to its maximum of 3.522. This is driven by a significant reallocation of representational power away from the low-frequency band (which drops to a minimum of 43.4%). Following this peak, the model enters a "specialization phase" (step 5000 onwards), where spectral entropy gradually decreases and frequency selectivity begins to recover, suggesting the model is refining and consolidating its newly learned complex representations.
Perhaps the most striking finding is the evolution of scale sensitivity, which increases in a two-step process. The first significant jump occurs between step 512 and step 1000, where sensitivity rises from 0.533 to 0.617. A second, larger leap happens between step 1000 and step 5000 (from 0.617 to 0.742), where it reaches its plateau. This shows that as the model’s spectral strategies become more complex, its representations become highly tuned to specific scales. Rather than developing a truly robust, scale-invariant understanding, the model specializes, making its learned positional strategies more brittle when the input scale is changed.
Our analysis is anchored by a very low reconstruction error across all checkpoints, validating the numerical stability of our wavelet methodology.
Training Step Spectral Entropy Frequency Selectivity Low Freq. Power (%) Scale Sens. (0.5x) Scale Sens. (0.25x)
0 2.277 0.757 77.0 0.523 0.539
128 2.291 0.760 76.0 0.522 0.539
512 3.169 0.369 54.4 0.533 0.549
1000 3.522 0.230 43.4 0.617 0.633
5000 3.381 0.294 48.6 0.742 0.762
10000 3.293 0.330 51.7 0.748 0.767
143000 3.010 0.454 59.7 0.740 0.756
Table 5: Evolution of spectral and scale invariance metrics for a Pythia model across training checkpoints.
@@ 6.1 Natural Evolution Toward Wavelet Behavior
RoPE’s rotational encoding introduces specific frequency components that propagate through the attention mechanism in a mathematically structured way. The rotation matrix $R(m\theta_{k})$ creates an inherent trade-off: larger $\theta$ provides precise positions but causes rapid rotation cycles that confuse distant relationships, while smaller $\theta$ better captures long-range patterns but blurs local positions. The wavelet-like properties we observe show how attention heads adapt to handle different frequency ranges created by these rotations.
@@ 7 Implications
Our findings have direct implications for the design, training, and specialization of language models. The discovery that attention heads spontaneously organize into frequency specialists suggests that models could be made more efficient by pre-initializing heads to a wavelet-like basis. This could accelerate convergence by bypassing the inefficient "exploratory phase" we identified during training. This functional understanding also provides a principled method for model pruning, creating smaller models by removing redundant frequency heads, and a powerful diagnostic tool for assessing training stability by monitoring spectral metrics.
Furthermore, this work provides a roadmap for model adaptation and future research. Understanding head specialization allows for targeted, task-specific fine-tuning, where one could adapt low-frequency heads for summarization or high-frequency heads for code analysis. This "wavelet framework" serves as a new lens for interpretability and establishes a new benchmark for positional encodings. The goal is no longer just to provide a position signal, but to find encodings that, like RoPE, act as a powerful inductive bias to catalyze the emergence of these optimal, multi-resolution strategies.
@@ 8 Conclusion
Our research demonstrates that Transformer models equipped with Rotary Position Embeddings (RoPE) are not merely subject to their theoretical limitations but actively overcome them by developing emergent, wavelet-like processing strategies. We have shown that this adaptation is not a minor artifact but a fundamental and multi-faceted characteristic of modern language models.
Indeed, we established that RoPE’s implicit relative encoding is uniquely instrumental in fostering the remarkable scale-invariance observed, distinguishing its behavior from models with absolute or explicit relative position biases. This adaptation is not static; rather, it is learned through distinct developmental phases. An initial "exploratory" stage, characterized by high spectral entropy, gives way to a "specialization" stage where models refine their strategies, becoming more efficient yet often more scale-dependent. Crucially, this dynamic learning process culminates in models that respect the fundamental Heisenberg-Gabor uncertainty principle, with larger and more advanced architectures demonstrating increasingly sophisticated and diversified strategies for managing this inherent trade-off between positional precision and spectral organization.
Looking forward, this "wavelet framework" offers a new lens for interpretability, allowing us to analyze what a model learns in terms of scale and frequency. Furthermore, understanding that these optimal structures are emergent rather than designed could inspire new architectural innovations, potentially leading to models that are either explicitly guided towards these solutions or are built with more efficient, innate multi-scale processing capabilities from the start.
@@ 9 Limitations
While our study provides strong evidence for the emergence of wavelet-like properties in RoPE-based models, its scope and interpretations have several limitations that open promising avenues for future research.
Our primary analysis, like much of the field, examines model properties at inference time after training is complete. While the Pythia checkpoint analysis offers a glimpse into the learning trajectory, a more fine-grained, continuous analysis of how these spectral properties evolve during the training process could reveal the precise dynamics and triggers for the observed phase shifts.
Although we analyzed a range of model sizes and architectures, our findings are based on a specific set of open-source language models trained predominantly on English text. Further research is needed to determine if these principles generalize across other modalities (e.g., vision, audio), data types (e.g., code, multilingual text), and proprietary architectures.
Broader impact
A potential societal implication of our work is that by demonstrating these beneficial properties intensify with model scale, our findings could be used to justify the trend of building ever-larger models. This risks exacerbating existing issues of computational and environmental resource concentration.
10 Acknowledgments
Research supported in part by European Union - Next Generation EU - namely by the MUR-PRIN 2022 project "2022REWNTE - Artificial Intelligence algorithms to track and detect Covid-19 vaccine-related infodemic on social media" - CUP no.B53D23020690006; in part by the projects FAIR under Grant PE0000013 and SERICS under Grant PE00000014 under the MUR National Recovery and Resilience Plan funded by the European Union-NextGenerationEU, and in part by the project Neural Reasoning over Open Data (NEREO) funded by the Italian Ministry of Education and Research (PRIN) under Grant 2022AEFHA, and in part by the project SEED funded by Sapienza University of Rome.